how to do 2/5X = -20
Answer by checkley79(3341)

2/5X = -20 CROSS MULTIPLY.
5X*-20=2
-100X=2
X=2/-100
X=.02 ANS.
OR
(2/5)X = -20
2X/5=-20
2X=5*-20
2X=-100
X=-100/2
X=-50 ANS.